OpenStudy (anonymous):

A rectangular room is 5 meters longer than it is wide, and its perimeter is 30 meters. Find the dimension of the room. The length is ? meters and the width is ? meters.

jimthompson5910 (jim_thompson5910):

P = 2L + 2W P = 2(W+5)+2W 30 = 2W + 10 + 2W 4W+10 = 30 Solve for W Then use this to find L = W+5

OpenStudy (anonymous):

w=5?

jimthompson5910 (jim_thompson5910):

good, so L is then ...

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

width = 5 meters length = 10 meters

OpenStudy (anonymous):

thank u..

jimthompson5910 (jim_thompson5910):

np